OpenStudy (anonymous):

ANYBODY?? In the below system, solve for y in the first equation. x + 3y = 6 2x - y = 10

OpenStudy (shamil98):

x = -3y + 6 2(-3y + 6) - y = 10 -6y + 12 - y =10 -7y = -2 y = 2/7 x + 3(2/7) = 6 x + 6/7 = 6 x = 5 (1/7)

OpenStudy (anonymous):

no you're wrong..://